# Existence results for Leibenson’s equation on Riemannian manifolds

### v1

@@ 3.1 Limiting procedure for solutions of the auxiliary problem
$\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\int_{0}^{\varepsilon}\int_{\Omega}u\varphi=\int_{\Omega}u(\cdot,0)\varphi.$
@@
$\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\int_{0}^{\varepsilon}\int_{\Omega}u\varphi=\int_{\Omega}u_{0}\varphi,$
which implies the initial condition since $\varphi\in C_{0}^{\infty}(\Omega)$ is arbitrary.
Lemma 3.14.
Let $u$ and $v$ be functions obtained in Lemma 3.10 with initial functions $u_{0}$ and $v_{0}$, respectively. Then, for any $t\in[0,T]$,
$\int_{\Omega}{\left(v-u\right)_{+}^{2}}(t)\leq\int_{\Omega}{\left(v_{0}-u_{0}\right)_{+}^{2}}.$ (3.54)
Proof. From Lemma 3.4, we have for any $t\in[0,T]$,
$\int_{\Omega}{\left(v_{N_{j}}-u_{N_{j}}\right)_{+}^{2}}(t)\leq\int_{\Omega}{\left(v_{0}-u_{0}\right)_{+}^{2}}.$
Sending $j\to\infty$ and using Lemma 3.10 we then obtain (3.54).
Corollary 3.15.
Assume also that $u_{0}\geq v_{0}$ a.e. in $\Omega$. Then $u\geq v$ a.e. in $\Omega\times(0,T)$.
@@ 4.1 Caccioppoli type inequalities
$\left[\int_{\Omega}(u-\theta_{1})_{+}^{\lambda}\eta^{p}\right]_{t_{1}}^{t_{2}}$ $+c_{1}\left(\frac{\theta_{1}}{\theta_{1}-\theta_{0}}\right)^{(q-1)(p-1)_{-}}\int_{Q}\left|\nabla\left((u-\theta_{1})_{+}^{\alpha}\eta\right)\right|^{p}$ (4.57)
@@
$|\nabla v|^{p}u^{(q-1)(p-1)}v^{\lambda-2}\eta^{p}$ $\geq\min\left(2^{1-p},2^{q(p-1)}\right)\alpha^{-p}\left|\nabla\left((u-\theta_{1})_{+}^{\alpha}\eta\right)\right|^{p}\left(\frac{\theta_{1}}{\theta_{1}-\theta_{0}}\right)^{(q-1)(p-1)_{-}}$
@@
$\left[\int_{\Omega}{v^{\lambda}\eta^{p}}\right]_{t_{1}}^{t_{2}}$ $+c_{1}\left(\frac{\theta_{1}}{\theta_{1}-\theta_{0}}\right)^{(q-1)(p-1)_{-}}\int_{Q}{\left|\nabla\left((u-\theta_{1})_{+}^{\alpha}\eta\right)\right|^{p}}$
@@
$\left[\int_{\Omega}{(u-\theta_{0})_{+}^{\lambda}\eta^{p}}\right]_{t_{1}}^{t_{2}}$ $+c_{1}\left(\frac{\theta_{1}}{\theta_{1}-\theta_{0}}\right)^{(q-1)(p-1)_{-}}\int_{Q}{\left|\nabla\left((u-\theta_{1})_{+}^{\alpha}\eta\right)\right|^{p}}$
@@ 4.2 Comparison in two cylinders
$J_{1}\leq\frac{Cr_{0}^{p}}{\left(\iota(B_{0})\mu(B_{0})\left(r_{0}-r_{1}\right)^{p}\right)^{\nu}(\theta_{1}-\theta_{0})^{\lambda\nu}\left(r_{0}-r_{1}\right)^{p}}\left(\frac{\theta_{1}}{\theta_{1}-\theta_{0}}\right)^{(q-1)(p-1)+\nu(q-1)(p-1)_{+}}J_{0}^{1+\nu},$ (4.71)
@@
$c_{1}\int_{t_{1}}^{T}\int_{B_{0}}\left|\nabla\left(v_{1}^{\alpha}\eta\right)\right|^{p}$ $\leq c_{2}\left(\frac{\theta_{1}}{\theta_{1}-\theta_{0}}\right)^{(q-1)(p-1)}\int_{t_{1}}^{T}\int_{B_{0}}v_{0}^{\sigma}|\nabla\eta|^{p}$
$\leq\frac{c_{3}}{\left(r_{0}-r_{1}\right)^{p}}\left(\frac{\theta_{1}}{\theta_{1}-\theta_{0}}\right)^{(q-1)(p-1)}J_{0},$ (4.73)
@@
$\quad\times\frac{c_{3}}{c_{1}\left(r_{0}-r_{1}\right)^{p}}\left(\frac{\theta_{1}}{\theta_{1}-\theta_{0}}\right)^{(q-1)(p-1)}J_{0},$
@@ 4.3 Iterations and the mean value theorem
Then, for the cylinder
$Q^{\prime}=\frac{1}{2}B\times[0,t],$
we have
@@
$\quad\times\left(\frac{\theta_{k+1}}{\theta_{k+1}-\theta_{k}}\right)^{(q-1)(p-1)+\nu(q-1)(p-1)_{+}}J_{k}^{1+\nu}.$
@@
$J_{k+1}$ $\leq$ $\frac{C2^{k\left(\lambda\nu+kp\left(1+\nu\right)+(q-1)(p-1)+\nu(q-1)(p-1)_{+}\right)}}{\left(\iota(B)\mu(B)R^{p}\right)^{\nu}\theta^{\lambda\nu}}J_{k}^{1+\nu}=\frac{A^{k}J_{k}^{1+\nu}}{\Theta}$
where $A=2^{\left(\lambda\nu+kp\left(1+\nu\right)+(q-1)(p-1)+\nu(q-1)(p-1)_{+}\right)}$ and $\Theta=c\left(\iota(B)\mu(B)\theta^{\lambda}R^{p}\right)^{\nu}$.
@@ 5 Global existence theorem
Assume that $pq\geq 1$ and that $u_{0}\in L^{1}(M)\cap L^{\infty}(M)$ is non-negative. Then there exists a non-negative bounded weak solution $u$ of the Cauchy problem
@@
Hence, after extending $u_{B}$ to $M$ by setting $u\equiv 0$ outside $B$, $u_{B}^{q}$ is a bounded sequence in $L^{p}((0,T),W^{1,p}(M))$. Thus, there is a subsequence $u_{B_{k}}^{q}$ and a function $w\in L^{p}((0,T),W^{1,p}(M))$ such that $u_{B_{k}}^{q}\to w$ weakly in $L^{p}((0,T),W^{1,p}(M))$ as $B_{k}\to M$. Since the sequence $\{u_{B}\}$ is monotone increasing in $B$ by Lemma 3.14, the limit is also a pointwise limit as $B_{k}\to M$. Thus, we obtain from (5.80) by sending $\lambda\to\infty$ that $u:=w^{1/q}$ is bounded. Using (5.80) with $\lambda=q+1$ and Fatou’s Lemma, we also obtain that $u\in L^{q+1}(M\times(0,T))$. Therefore, we can argue as in Lemma 3.12 and get that $u\in C([0,T],L^{q+1}(M))$.
@@
for any $\psi\in W^{1,1+\frac{1}{q}}\left((0,T);L^{1+\frac{1}{q}}(M)\right)\cap L^{p}\left((0,T);W^{1,p}(M)\right)$ and that $u(\cdot,0)=u_{0}$. If $u$ and $v$ are limit functions obtained in this way with initial function $u_{0}$ and $v_{0}$, respectively, we have by Lemma 3.14, for $t\in[0,T]$,
$\int_{M}{\left(u_{B_{k}}-v_{B_{k}}\right)_{+}^{2}}(t)\leq\int_{M}{\left(u_{0}-v_{0}\right)_{+}^{2}}.$
Thus, for $k\to\infty$, we get
$\int_{M}{\left(u-v\right)_{+}^{2}}(t)\leq\int_{M}{\left(u_{0}-v_{0}\right)_{+}^{2}}.$ (5.82)
From the estimate (5.82) we get a solution in $M\times(0,\infty)$, which finishes the proof.

### v2

@@ 3.1 Limiting procedure for solutions of the auxiliary problem
$\frac{1}{\varepsilon}\int_{0}^{\varepsilon}\int_{\Omega}u\varphi\to\int_{\Omega}u(\cdot,0)\varphi\quad\textnormal{as}~\varepsilon\to 0.$
@@
$\frac{1}{\varepsilon}\int_{0}^{\varepsilon}\int_{\Omega}u\varphi\to\int_{\Omega}u_{0}\varphi\quad\textnormal{for}~\varepsilon\to 0$
, which implies the initial condition since $\varphi\in C_{0}^{\infty}(\Omega)$ is arbitrary.
@@ 4.1 Caccioppoli type inequalities
$\left[\int_{\Omega}(u-\theta_{1})_{+}^{\lambda}\eta^{p}\right]_{t_{1}}^{t_{2}}$ $+c_{1}\left(\frac{\theta_{1}}{\theta_{1}-\theta_{0}}\right)^{-(q-1)(p-1)_{-}}\int_{Q}\left|\nabla\left((u-\theta_{1})_{+}^{\alpha}\eta\right)\right|^{p}$
@@
$|\nabla v|^{p}u^{(q-1)(p-1)}v^{\lambda-2}\eta^{p}$ $\geq\min\left(2^{1-p},2^{q(p-1)}\right)\alpha^{-p}\left|\nabla\left((u-\theta_{1})_{+}^{\alpha}\eta\right)\right|^{p}\left(\frac{\theta_{1}}{\theta_{1}-\theta_{0}}\right)^{-(q-1)(p-1)_{-}}$
@@
$\left[\int_{\Omega}{v^{\lambda}\eta^{p}}\right]_{t_{1}}^{t_{2}}$ $+c_{1}\left(\frac{\theta_{1}}{\theta_{1}-\theta_{0}}\right)^{-(q-1)(p-1)_{-}}\int_{Q}{\left|\nabla\left((u-\theta_{1})_{+}^{\alpha}\eta\right)\right|^{p}}$
@@
$\left[\int_{\Omega}{(u-\theta_{0})_{+}^{\lambda}\eta^{p}}\right]_{t_{1}}^{t_{2}}$ $+c_{1}\left(\frac{\theta_{1}}{\theta_{1}-\theta_{0}}\right)^{-(q-1)(p-1)_{-}}\int_{Q}{\left|\nabla\left((u-\theta_{1})_{+}^{\alpha}\eta\right)\right|^{p}}$
@@ 4.2 Comparison in two cylinders
$J_{1}\leq\frac{Cr_{0}^{p}}{\left(\iota(B_{0})\mu(B_{0})\left(r_{0}-r_{1}\right)^{p}\right)^{\nu}(\theta_{1}-\theta_{0})^{\lambda\nu}\left(r_{0}-r_{1}\right)^{p}}\left(\frac{\theta_{1}}{\theta_{1}-\theta_{0}}\right)^{|q-1|(p-1)+\nu(q-1)(p-1)_{+}}J_{0}^{1+\nu},$ (4.78)
@@
$c_{1}\int_{t_{1}}^{T}\int_{B_{0}}\left|\nabla\left(v_{1}^{\alpha}\eta\right)\right|^{p}$ $\leq c_{2}\left(\frac{\theta_{1}}{\theta_{1}-\theta_{0}}\right)^{|q-1|(p-1)}\int_{t_{1}}^{T}\int_{B_{0}}v_{0}^{\sigma}|\nabla\eta|^{p}$
$\leq\frac{c_{3}}{\left(r_{0}-r_{1}\right)^{p}}\left(\frac{\theta_{1}}{\theta_{1}-\theta_{0}}\right)^{|q-1|(p-1)}J_{0},$ (4.80)
@@
$\quad\times\frac{c_{3}}{c_{1}\left(r_{0}-r_{1}\right)^{p}}\left(\frac{\theta_{1}}{\theta_{1}-\theta_{0}}\right)^{|q-1|(p-1)}J_{0},$
@@ 4.3 Iterations and the mean value theorem
Then, for the cylinder $Q^{\prime}=\frac{1}{2}B\times[0,t],$ we have
@@
$\quad\times\left(\frac{\theta_{k+1}}{\theta_{k+1}-\theta_{k}}\right)^{|q-1|(p-1)+\nu(q-1)(p-1)_{+}}J_{k}^{1+\nu}.$
@@
$J_{k+1}$ $\leq$ $\frac{C2^{k\left(\lambda\nu+p\left(1+\nu\right)+|q-1|(p-1)+\nu(q-1)(p-1)_{+}\right)}}{\left(\iota(B)\mu(B)R^{p}\right)^{\nu}\theta^{\lambda\nu}}J_{k}^{1+\nu}=\frac{A^{k}J_{k}^{1+\nu}}{\Theta}$
where $A=2^{\left(\lambda\nu+p\left(1+\nu\right)+|q-1|(p-1)+\nu(q-1)(p-1)_{+}\right)}\geq 1$ and $\Theta=c\left(\iota(B)\mu(B)\theta^{\lambda}R^{p}\right)^{\nu}$.
@@ 5 Global existence theorem
Assume that $pq\geq 1$ and that $u_{0}\in L^{1}(M)\cap L^{\infty}(M)$ is non-negative. Then there exists a unique non-negative bounded weak solution $u$ of the Cauchy problem
@@
Hence, after extending $u_{B}$ to $M$ by setting $u\equiv 0$ outside $B$, $u_{B}^{q}$ is a bounded sequence in $L^{p}((0,T),W^{1,p}(M))$. Thus, there is a subsequence $u_{B_{k}}^{q}$ and a function $w\in L^{p}((0,T),W^{1,p}(M))$ such that $u_{B_{k}}^{q}\to w$ weakly in $L^{p}((0,T),W^{1,p}(M))$ as $B_{k}\to M$. Since the sequence $\{u_{B}\}$ is monotone increasing in $B$ by (2.34), the limit is also a pointwise limit as $B_{k}\to M$. Thus, we obtain from (5.87) by sending $\lambda\to\infty$ that $u:=w^{1/q}$ is bounded. Using (5.87) with $\lambda=q+1$ and Fatou’s Lemma, we also obtain that $u\in L^{q+1}(M\times(0,T))$. Therefore, we can argue as in Lemma 3.9 and get that $u\in C([0,T],L^{q+1}(M))$.
@@
for any $\psi\in W^{1,1+\frac{1}{q}}\left((0,T);L^{1+\frac{1}{q}}(M)\right)\cap L^{p}\left((0,T);W^{1,p}(M)\right)$ and that $u(\cdot,0)=u_{0}$.
From Lemma 2.5 we get a unique solution in $M\times(0,\infty)$, which finishes the proof.